OpenStudy (anonymous):

If the standard for is (x-13)^2 +(y-9)^2=121 the center of the circle is (15/2,9) the what is the radius of the circle?

OpenStudy (shinalcantara):

please check the given.. 'cause if that's the equation, then you'll have (13,9) as the center of the circle..

OpenStudy (shinalcantara):

the standard form for the equation of a circle is: \[(x-h)^{2} + (y-k)^{2} = r ^{2}\] where (h,k) is the center

OpenStudy (anonymous):

on my homework i put (13,9)..heres the peoblem

OpenStudy (shinalcantara):

by inspection, you'll find the radius as

OpenStudy (shinalcantara):

since you already have (h,k) and r values, substitute them to the standard equation..

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Do you know the standard form?

OpenStudy (shinalcantara):

by substituting the values you'll have: \[(x - \frac{ 15 }{ 2 })^{2} + (y - 9)^{2} = 5.5^{2} \] since 5.5^2 is 30.25, you may have it as 30.25

OpenStudy (anonymous):

ahhh I see, ok thank you lots